OpenStudy (lilai3):

Plz help! It will be greatly appreciated; WILL AWARD MEDAL!!! Find the surface area of a cone: base area, 805 mm^2; slant height, 30 mm. Surface area?

OpenStudy (agent0smith):

Area of base is \[A = πr^2\] So knowing that the base area is 805mm^2, we can find r. Then we can find total SA from: \[SA = πr^2 + πrl\] where l is the slant height.

OpenStudy (agent0smith):

find r from\[ 805= \pi r^2\]divide both sides by pi\[\frac{ 805 }{\pi }= r^2\]take sqrt of both sides \[\sqrt { \frac{ 805 }{\pi } }= r\] Now use \[SA = πr^2 + πrl \]

OpenStudy (lilai3):

u just answereed my question(:

OpenStudy (lilai3):

thank you @agent0smith . can i fan you too? err... can you fan me too cuz we can chat and stuff and i can help you and stuff...? well it's okay to reject my offer. lOLS

OpenStudy (agent0smith):

lol you can chat with me even if i don't fan you :)

OpenStudy (lilai3):

oh really good. let's experiment

OpenStudy (lilai3):

wait... but you only showed me the r you didn't finish yet

OpenStudy (lilai3):

we're finding the surface area of the whole cone remebmer?

OpenStudy (agent0smith):

Yeah... but i wanted you to try that part :P you can find r from what i showed. Then find surface area from the formula i gave.

OpenStudy (lilai3):

well can you show me? cuz i have a test tomorro and i suck at this. show me please?

OpenStudy (agent0smith):

\[r = \sqrt { \frac{ 805 }{\pi } }\]find this from your calculator. Then, put that number into this formula as r, along with 30mm for l \[SA = π r^2 + π r l\] and you already know the base area is 805 mm^2 so you can make it a bit simpler: \[SA = 805 + π r l\] Now all you do is plug in r and l.

OpenStudy (agent0smith):

Oh and you can put in 30 for l, now all you need to do is enter r\[SA = 805 +30 π r \]